nesha:

A sequence is defined by the recursive formula f(n + 1) = 1.5f(n). Which sequence could be generated using the formula?

TheSmartOne:

Hi there, welcome to QuestionCove! Let's plug in some numbers as that's the best way to get to the answer! If we say f(0)= 1 What do you get when you plug in n=0 What do you get when n = 1, 2, 3 Do you notice a pattern between the numbers? Now look at your answer choices and see which one has that same pattern

nesha:

–12, –18, –27, ... –20, 30, –45, ... –18, –16.5, –15, ... –16, –17.5, –19, ...

TheSmartOne:

Thank you for providing the answer choices. Let's say that f(0) = 1 now plug in n = 0 f(0+1) = 1.5*f(0) f(1) = 1.5 * f(0) Plug in 1 for f(0) because that's what I said it was earlier. I just made up a value because it's important that we understand the pattern/relationship between the numbers f(1) = 1.5*1 f(1) = 1.5 So now can you find f(2) is? You have to plug in n= 1

TheSmartOne:

Once you figure out the pattern, we can look to see which answer choice has the same correct pattern.

TheSmartOne:

There are multiple correct answers because they didn't tell us what one value would be when you have a certain n. So that's why we need to find the pattern

TheSmartOne:

But among your answer choices, they only listed one possible correct answer choice. Please do ask any questions you have if you're confused